# Memory effects in a dynamical decoupling process by parity kicks

### v1

@@ Abstract
We establish a simple quantitative relationship between the environmental memory effects and the characteristics in a dynamical decoupling process. In contrast to previous works, our measures of non-Markovianity are tailored and extended to evaluate the strength of memory effects in dynamical decoupling. We find that if each kick commutes with the dynamical map of the uncontrolled system, then the change of the final dynamical map or the final state brought by the control (called the “effect of control”) is upper (lower) bounded by the summation (difference) of the strengths of memory effects with and without control. Our finding is illustrated with a dissipative quantum Rabi model by numerical simulations where one or many cycles of parity kicks are implemented on the qubit. Besides, the results show that under certain conditions, the effect of control or the increase of performance by the control may be simply proportional to the strength of memory effects with or without control.
@@ I Introduction
It is well known that the non-Markovian character of the environment (memory effects) is crucial for the success of dynamical decoupling since the period of the kicks has to be small compared with the fastest time scale of the interactions with the environment [1, 3, 2, 4]. From another perspective, in the time scale of the period of the kicks, the information lost from the system has to be preserved in the environment and exploited to influence the system’s future dynamics. Therefore, the system evolution under control depends on its history which is the very essence of the non-Markovian character. In recent years, the interplay between dynamical decoupling and the environmental memory effects has received much attention [19, 20, 21, 22, 23, 24, 25, 26, 27]. On one hand, the investigation of this issue is beneficial for the advance of quantum information processing. On the other hand, it sheds light on the concept of quantum non-Markovianity [28, 29, 30, 31] which is of fundamental interest.
@@
Particularly, the quantitative relations between the characteristics of dynamical decoupling and strength of environmental memory effects (non-Markovianity) have been studied in several works, yet not so extensively. For example, the efficiency of dynamical decoupling is compared with the Breure-Laine-Pilo measure of non-Markovianity [34] in uncontrolled or controlled dynamics in a dephsaing model in Refs. [19] and [24]. This measure [34] is defined as the total increments of distinguishability in time evolutions maximized over all the possible pair of initial states and associated with the information backflow. However, their results indicate that the dependence of the control efficiency on the non-Markovianity is not quite straightforward [19, 24] (neither proportional nor monotonic). Recently, using quantum resource theories [32, 33], the authors in Ref. [25] show that quantum non-Markovianity, interpreted as the multitime correlations, is a resource for dynamical decoupling. The theory is formulated in the framework of process tensor whose characterization scales exponentially with the number of kicks. In their work, the kicks are assumed to be free transformations that do not increase the non-Markovianity and the success of dynamical decoupling is quantified by the total temporal correlations in the process tensor. Before and after a coarse-graining process to a less number of interventions [25], it is seen that the increase of the total temporal correlations in the process tensor is accompanied by the consumption of non-Markovian correlations.
@@
Given the conceptual and computational complexities in Ref. [25], the aim of this paper is to establish a simple relation that quantitatively links the dynamical decoupling characteristics and the strength of memory effects. To accomplish this, we adopt and extend the measures of non-Markovianity proposed in our previous works [35, 36] to quantify the strength of memory effects in a dynamical decoupling process (or that without control). Originally, our non-Markovianity measures are defined as the maximal difference of completely positive dynamical maps $T(t_{2},t_{0})$ and $T(t_{2},t_{1})T(t_{1},t_{0})$ [35] (or those acting on a system initial state [36]) by optimizing over $t_{1}$ and $t_{2}$ ($t_{0}$ is fixed). In this work, we fix $t_{1}$ and $t_{2}$ to quantify the memory effects in a dynamical decoupling process with two kicks at $t_{1}$ and $t_{2}$ and further extend our measures to their multitime forms for more kicks. Compared with the measures used in Refs. [19, 24] or [25], the measures in this work have the following merits. First, they can be directly interpreted as the future-past dependence of the system’s dynamics, which is the essence of non-Markovianty. Second, the calculation of memory effects is easy since the description of the dynamical maps [35] or quantum states [36] in our measures does not scale with the number of kicks and requires no optimizations over the initial states [34] or time instants [35, 36]. Last, our measures generally give non-zero strength of memory effects for any open quantum process without the Born-Markov approximation, even that without information backflow [34], e.g., a near-Markovnian dynamics [35, 36] or Zeno dynamics [36].
@@
Intuitively, the dynamical decoupling of a near-Markovian quantum process leads to a low efficiency and very weak memory effects, meanwhile, a perfectly decoupled dynamics may become effectively unitary and thus Markovian, again leading to very weak memory effects. Thus there may not exist a monotonic relation between the dynamical decoupling efficiency and the strength of memory effects in the controlled process. Instead, we focus on the change of the system’s final dynamical map (final state) brought by dynamical decoupling, which we called the “effect of control”, and the strengths of memory effects in both the controlled and uncontrolled processes. Our main finding is that if each kick commutes with the dynamical map of the uncontrolled system, then the effect of control is upper (lower) bounded by the summation (difference) of the strengths of memory effects in the controlled and uncontrolled processes. That is, the non-Markovianity of the controlled and uncontrolled process together contribute to the change of the final dynamical map (final state) in general. We illustrate our finding by numerically simulating a dissipative quantum Rabi model where parity kicks [2] are applied to the qubit to eliminate the influence of its environment (an oscillator in a Markovian thermal reservoir). Results for one and many cycles of parity kicks in different regimes coincide with our theory. Moreover, we find that, in many cases, the effect of control may be approximately equal to the increase of performance brought by the control and proportional to the strength of memory effects in the controlled or uncontrolled process.
@@
The paper is organized as follows: In Sec. II we briefly introduce the theory of parity kicks and the standard dynamical decoupling. In Sec. III, our non-Markovianity measures are reviewed, tailored and extended to evaluate the memory effects in dynamical decoupling. Next, these measures are applied to a dynamical decoupling process with two or more kicks in Sec. IV to establish quantitative connections between the effect of control and the strengths of memory effects. In Sec.V, we illustrate our finding with a dissipative quantum Rabi model controlled by one and many cycles of parity kicks in different regimes. At last, we summarize our work in Sec. VI.
@@ II Dynamical Decoupling
[…]
@@ II.2 General case
$K_{j}=\left\{\begin{array}[]{rcl}g_{j+1}g^{{\dagger}}_{j}&&(j=1,2,\cdots n_{g}-1)\\ g^{{\dagger}}_{n_{g}}&&(j=n_{g})\\ \end{array}\right.$ (14)
with $n_{g}=m$. Using the above construction, Eq. (7) can be understood by $K_{1}=g_{2}g_{1}^{{\dagger}}=K_{P}$ and $K_{2}=g^{{\dagger}}_{2}=K_{P}$ where $g_{1}=I$ and $g_{2}=K_{P}$ are the elements of a group of order 2. Using Eq. (14), it is easy to see that the condition
@@ III.1 Measure of memory effects
The measure of non-Markovianity defined in Ref. [35] deals with a quantum process governed by a system-environment Hamiltonian (possibly time-dependent) $H(t)=H_{S}(t)+H_{E}(t)+H_{SE}(t)$ like Eq. (1) or Eq. (3), an arbitrary system initial state $\rho_{S}(t_{I})$ and a fixed initial state of the environment $\rho_{E}(t_{I})$. Here $t_{I}$ is an arbitrary initial time of an evolution conditioned by
@@
i.e., the system is uncorrelated with the environment before an evolution. In general, the initial state of the environment is given by [31]
[…]
If the environment Hamiltonian $H_{E}$ is time-independent and the environmental initial state is a steady state of $H_{E}$ (e.g., a thermal state), Eq. (16) reduces to $\rho_{SE}(t_{I})=\rho_{S}(t_{I})\otimes\rho_{E}$.
We define $T(t_{b},t_{a})$ as the dynamical map that transfers $\rho_{S}(t_{a})$ to $\rho_{S}(t_{b})$ ($t_{a}\leqslant t_{b}$) where $t_{a}$ is the initial time of the evolution ($t_{a}=t_{I}$) [35], i.e.,
@@
$=$ $\text{Tr}_{E}[U(t_{b},t_{a})\rho_{S}(t_{a})\otimes\rho_{E}(t_{a})U(t_{b},t_{a})^{{\dagger}}].$
[…]
Let $t_{0}\leqslant t_{1}\leqslant t_{2}$, a (time-dependent) quantum process is Markovian if the dynamical maps (denoted by $\Lambda_{M}$) satisfy the divisibility condition
$\Lambda_{M}(t_{2},t_{0})=\Lambda_{M}(t_{2},t_{1})\Lambda_{M}(t_{1},t_{0}),$ (19)
where each dynamical map is uniquely defined and a UDM. Remark that the dynamical map $\Lambda_{M}(t_{2},t_{1})$ is a UDM if and only if it is induced by
$\rho_{S}(t_{2})$ $=$ $\Lambda_{M}(t_{2},t_{1})\rho_{S}(t_{1})$ (20)
$=$ $\text{Tr}_{E}[U(t_{2},t_{1})\rho_{S}(t_{1})\otimes\rho_{E}(t_{1})U(t_{2},t_{1})^{{\dagger}}],$
where $\rho_{E}(t_{1})$ is fixed and $\rho_{S}(t_{1})$ is arbitrary [40]. It is observed that the Markovian dynamical map $\Lambda_{M}(t_{2},t_{1})$ is unique no matter $t_{I}<t_{1}$ or $t_{I}=t_{1}$. Therefore, $\Lambda_{M}(t_{2},t_{1})=T(t_{2},t_{1})$ according to our definition. Then, the Markovian divisibility condition (19) can be expressed by
@@
In contrast to the non-Markovianity criteria in Refs. [43, 44], the violation of Eq. (21) is manifested by the inequality
@@
where all the dynamical maps are TPCP.
The physical meaning of Eq. (22) can be interpreted as memory effects in two ways. This is done by acting the left-hand and the right-hand sides on a system initial state $\rho_{S}(t_{0})$ and checking their final states at $t_{1}$ as illustrated in Fig. 1. We call the evolutions corresponding to $\rho_{S}(t_{2})=T(t_{2},t_{0})\rho_{S}(t_{0})$, $\rho_{S}(t_{1})=T(t_{1},t_{0})\rho_{S}(t_{0})$ and $\rho^{\prime}_{S}(t_{2})=T(t_{2},t_{1})\rho_{S}(t_{1})$ evolution A, B and C, respectively. If the inequality holds, there exist initial states $\rho_{S}(t_{0})$ such that $\rho_{S}(t_{2})\neq\rho^{\prime}_{S}(t_{2})$. From the perspective of the system, the inequality could be interpreted as the future-past dependence of the system’s evolution. At the moment $t_{1}$, evolution A and C have the same system state $\rho_{S}(t_{1})$ but different histories. Evolution A has a history in time interval $[t_{0},t_{1}]$ encoded in $\rho_{SE}(t_{1})=U(t_{1},t_{0})\rho_{S}(t_{0})\otimes\rho_{E}(t_{0})U(t_{1},t_{0})^{{\dagger}}$. In contrast, evolution C has no history before $t_{1}$ since the initial condition of evolution C is $\rho_{S}(t_{1})\otimes\rho_{E}(t_{1})$ with fixed $\rho_{E}(t_{1})$ defined by Eq. (17). Therefore, $\rho_{S}(t_{2})\neq\rho^{\prime}_{S}(t_{2})$ is an evidence that the future state (after $t_{1}$) of the system depends on its history (in $[t_{0},t_{1}]$) in this quantum process. From the perspective of the environment, the inequality could be understood from the change of the environment state between evolution B and C and its influence on the future system. At the end of evolution B, the system’s history in $[t_{0},t_{1}]$ has been encoded in $\rho_{SE}(t_{1})$ as in evolution A. Then, at the beginning of evolution C, the history information is erased by $T(t_{2},t_{1})$ such that
$\rho_{SE}(t_{1})\rightarrow\rho_{S}(t_{1})\otimes\rho_{E}(t_{1})$ (23)
where $\rho_{E}(t_{1})$ is independent of the system. In contrast, such an erasure process never happens in evolution A. Therefore, $\rho_{S}(t_{2})\neq\rho^{\prime}_{S}(t_{2})$ signifies that the environment (as well as the correlations between the system and environment [40]) remembers the system’s history information which in turn influences the system’s future.
Figure 1: Physical interpretation of Eq. (22) as memory effects. From the system perspective, the system states in evolution A and C are the same at $t_{1}$. However, the system has different evolution histories in evolution A and C. From the environment perspective, the environment in evolution B is initialized by $T(t_{2},t_{1})$ at $t_{1}$ so that the system’s history in evolution B is erased. Therefore, $\rho_{S}(t_{2})\neq\rho^{\prime}_{S}(t_{2})$ demonstrates that the system’s history in $[t_{0},t_{1}]$ have an influence on the system’s future at $t_{2}$.
Consider a quantum process where all the interested evolutions happens in the interval $[t_{0},t_{max}]$. Let $t_{0}\leqslant t_{1}\leqslant t_{2}\leqslant t_{max}$ and $t_{0}=0$ by default. In Ref. [35], the non-Markovnianity of a quantum process is defined as the maximal trace distance $D$ between the Choi-Jamiółkowski matrices [41, 42] of $T(t_{2},t_{0})$ and $T(t_{2},t_{1})T(t_{1},t_{0})$ while optimizing over $t_{1}$ and $t_{2}$, specifically,
@@
Here $D(\rho_{1},\rho_{2})=\frac{1}{2}\|\rho_{1}-\rho_{2}\|$ is the trace distance between two density matrices where $\|A\|=\text{Tr}(\sqrt{A^{\dagger}A})$ is the trace norm of an operator $A$. The Choi-Jamiółkowski matrix of a dynamical map $\Lambda$ is given by $\rho_{\Lambda}=(\mathbb{I}\otimes\Lambda)(|\psi\rangle\langle\psi|)$ where $|\psi\rangle=\frac{1}{\sqrt{d}}\sum_{i=1}^{d}|i\rangle|i\rangle$ is a maximally entangled state of an ancillary system and the system of interest (both $d$-dimensional) and $\mathbb{I}$ is the identity map in the space of the ancillary system. To fully characterize the non-Markovian features of a quantum process, one can let $t_{max}=\infty$ [35] so that all the possible $t_{1}$ and $t_{2}$ are taken into account. Using the dynamical maps $T(t_{b},t_{a})$ in the definition Eq.(24), the measure does not depend on the system’s initial state.
To investigate the influence of particular initial states on the memory effects, we recently define
@@
as the strength of memory effects conditioned on the initial state $\rho_{S}(t_{0})$ [36]. Note that a nonzero $N_{M}[\rho_{S}(t_{0})]$ is a sufficient condition for the non-Markovianity of a quantum process defined by Eq.(22) or Eq.(24). Compared with Eq.(24), the initial-state-dependent measure of memory effects is easier to calculate and verify. This approach is used to study the early-time superradiance where a finite $t_{max}$ is adopted [36]. In this work, the two measures mentioned above are both considered, which allows us to deal with the problems of controlling dynamical maps and quantum states (with any initial state) by dynamical decoupling.

### v2

@@ Abstract
We establish a simple quantitative relationship between the environmental memory effects and the characteristics in a dynamical decoupling process. In contrast to previous works, our measures of non-Markovianity are tailored and extended to evaluate the strength of memory effects in dynamical decoupling. We find that if each kick commutes with the dynamical map of the uncontrolled system, then the change of the final dynamical map or the final state brought by the control (called the “effect of control”) is upper (lower) bounded by the summation (difference) of the strengths of memory effects with and without control. We propose sufficient conditions for the commutation relation for parity kicks and illustrate our finding with a dissipative quantum Rabi model by numerical simulations where one or many cycles of parity kicks are implemented on the qubit. Besides, the results show that under certain conditions, the effect of control or the increase of performance by the control may be simply proportional to the strength of memory effects with or without control.
@@ I Introduction
It is well known that the non-Markovian character of the environment (memory effects) is crucial for the success of dynamical decoupling since the period of the kicks has to be small compared with the fastest time scale of the interactions with the environment [1, 3, 2, 4]. From an informational perspective, in the time scale of the period of the kicks, the information lost from the system has to be preserved in the environment and exploited to influence the system’s future dynamics. Therefore, the system evolution under control depends on its history which is the very essence of the non-Markovian character. In recent years, the interplay between dynamical decoupling and the environmental memory effects has received much attention [19, 20, 21, 22, 23, 24, 25, 26, 27]. On one hand, the investigation of this issue is beneficial for the advance of quantum information processing. On the other hand, it sheds light on the concept of quantum non-Markovianity [28, 29, 30, 31] which is of fundamental interest.
@@
Particularly, the quantitative relations between the characteristics of dynamical decoupling and strength of environmental memory effects (non-Markovianity) have been studied in several works, yet not so extensively. For example, the efficiency of dynamical decoupling is compared with the Breuer-Laine-Pilo measure of non-Markovianity [32] in uncontrolled or controlled dynamics in a dephsaing model in Refs. [19] and [24]. This measure [32] is defined as the total increments of distinguishability in time evolutions maximized over all the possible pair of initial states and associated with the information backflow. However, their results indicate that the dependence of the control efficiency on the non-Markovianity is not quite straightforward [19, 24] (neither proportional nor monotonic). Recently, using quantum resource theories [33, 34], the authors in Ref. [25] show that quantum non-Markovianity, interpreted as the multitime correlations, is a resource for dynamical decoupling. The theory is formulated in the framework of process tensor whose characterization scales exponentially with the number of kicks. In their work, the kicks are assumed to be free transformations that do not increase the non-Markovianity and the success of dynamical decoupling is quantified by the total temporal correlations in the process tensor. Before and after a coarse-graining process to a smaller number of interventions [25], it is seen that the increase of the total temporal correlations in the process tensor is accompanied by the consumption of non-Markovian correlations.
@@
Given the conceptual and computational complexities in Ref. [25], the aim of this paper is to establish a simple relation that quantitatively links the dynamical decoupling characteristics and the strength of memory effects. To accomplish this, we adopt and extend the measures of non-Markovianity proposed in our previous works [35, 36] to quantify the strength of memory effects in a dynamical decoupling process (or that without control). Originally, our non-Markovianity measures are defined as the maximal difference of completely positive dynamical maps $T(t_{2},t_{0})$ and $T(t_{2},t_{1})T(t_{1},t_{0})$ [35] (or those acting on a system initial state [36]) by optimizing over $t_{1}$ and $t_{2}$ ($t_{0}$ is fixed). In this work, we fix $t_{1}$ and $t_{2}$ to quantify the memory effects in a dynamical decoupling process with two kicks at $t_{1}$ and $t_{2}$ and further extend our measures to their multitime forms for more kicks. Compared with the measures used in Refs. [19, 24] or [25], the measures in this work have the following merits. First, they can be directly interpreted as the past-future dependence, i.e., the dependence of the system’s future dynamics on its history, which is the essence of non-Markovianity. Second, the calculation of memory effects is computationally efficient since the description of the dynamical maps [35] or quantum states [36] in our measures does not scale with the number of kicks and requires no optimizations over the initial states [32] or time instants [35, 36]. Last, our measures reflect the quality of the Born-Markov approximation for a quantum process [35, 36].
@@
Intuitively, the dynamical decoupling of a near-Markovian quantum process leads to a bad performance (e.g., low fidelity between the final state and the target state.) and very weak memory effects, meanwhile, a perfectly decoupled dynamics may become effectively unitary and thus Markovian, again leading to very weak memory effects. Thus there may not exist a monotonic relation between the dynamical decoupling performance and the strength of memory effects in the controlled process. Instead, we focus on the change of the system’s final dynamical map (final state) brought by dynamical decoupling, which we called the “effect of control”, and the strengths of memory effects in both the controlled and uncontrolled processes. Our main finding is that if each kick commutes with the dynamical map of the uncontrolled system, then the effect of control is upper (lower) bounded by the summation (difference) of the strengths of memory effects in the controlled and uncontrolled processes. That is, the non-Markovianity of the controlled and uncontrolled process together contribute to the change of the final dynamical map (final state) in general. We propose a class of models where a parity kick [2] commutes with the dynamical map without control. After that, we illustrate our finding by numerically simulating a dissipative quantum Rabi model where parity kicks are applied to the qubit to eliminate the influence of its environment (an oscillator in a Markovian thermal reservoir). Results for one and many cycles of parity kicks in different regimes coincide with our theory. Moreover, we find that, in many cases, the effect of control may be approximately equal to the increase of performance brought by the control and proportional to the strength of memory effects in the controlled or uncontrolled process.
@@
The paper is organized as follows: In Sec. II we briefly introduce the theory of parity kicks and the standard dynamical decoupling. In Sec. III, our non-Markovianity measures are reviewed, tailored and extended to evaluate the memory effects in dynamical decoupling. Next, these measures are applied to a dynamical decoupling process with two or more kicks in Sec. IV to establish quantitative connections between the effect of control and the strengths of memory effects. In Sec.V, we propose sufficient conditions for the commutation between a parity kick and the uncontrolled dynamical map and illustrate our finding with a dissipative quantum Rabi model in different regimes. At last, we summarize our work in Sec. VI.
@@ II Dynamical Decoupling
and the system-environment evolution under control is determined by $U(t)=\mathcal{T}e^{-i\int_{0}^{t}H_{c}(\tau)d\tau}$ where $\mathcal{T}$ is the time-ordering operator ($\hbar=1$ throughout the paper).
@@ II.2 General case
[…]
with $n_{g}=m$. Using the above construction, Eq. (7) can be understood by $K_{1}=g_{2}g_{1}^{{\dagger}}=K_{P}$ and $K_{2}=g^{{\dagger}}_{2}=K_{P}$ where $g_{1}=I$ and $g_{2}=K_{P}$ are the elements of a group of order 2. Thus the parity kick scheme can be cast into the framework of dynamical decoupling. Using Eq. (14), it is easy to see that the condition
@@ III.1 Measure of memory effects
The measure of non-Markovianity defined in Ref. [35] deals with a quantum process governed by a system-environment Hamiltonian (possibly time-dependent) $H(t)=H_{S}(t)+H_{E}(t)+H_{S\!E}(t)$ like Eq. (1) or Eq. (3), an arbitrary system initial state $\rho_{S}(t_{I})$, a fixed initial state of the environment $\rho_{E}(t_{I})$, and the condition
@@
where $t_{I}$ is an arbitrary initial time. In general, the initial state of the environment is given by [31]
$\rho_{E}(t_{I})=\mathcal{T}e^{-i\int_{0}^{t_{I}}H_{E}(t^{\prime})dt^{\prime}}\rho_{E}(0).$ (17)
We define the universal dynamical map (UDM) [39] $T(t_{b},t_{a})$ as the dynamical map that transfers any $\rho_{S}(t_{a})$ to $\rho_{S}(t_{b})$ ($t_{a}\leqslant t_{b}$) [35], i.e.,
@@
$=$ $\mathrm{Tr}_{E}[U(t_{b},t_{a})\rho_{S}(t_{a})\otimes\rho_{E}(t_{a})U(t_{b},t_{a})^{{\dagger}}].$
where $t_{a}=t_{I}$ and $U(t_{b},t_{a})=\mathcal{T}e^{-i\int_{t_{a}}^{t_{b}}H(\tau)d\tau}$. Let $t_{0}\leqslant t_{1}\leqslant t_{2}$, then the Markovian divisibility condition can be expresssed as [35, 39]
@@
In contrast to the non-Markovianity criteria in Refs. [40, 41], the violation the Eq. (19) is manifested by the inequality
@@
By comparing the evolutions $\rho_{S}(t_{2})=T(t_{2},t_{0})\rho_{S}(t_{0})$ and $\rho^{\prime}_{S}(t_{2})=T(t_{2},t_{1})\rho_{S}(t_{1})$ where the system states at $t_{1}$ are both $\rho_{S}(t_{1})=T(t_{1},t_{0})\rho_{S}(t_{0})$, the fact $\rho_{S}(t_{2})\neq\rho^{\prime}_{S}(t_{2})$ can be interpreted as the dependence of the system’s future dynamics (from $t_{1}$ to $t_{2}$) on its history (from $t_{0}$ to $t_{1}$). On the other hand, in the process $\rho^{\prime}_{S}(t_{2})=T(t_{2},t_{1})T(t_{1},t_{0})\rho_{S}(t_{0})$, the system’s history (from $t_{0}$ to $t_{1}$) encoded in $\rho_{S\!E}(t_{1})$ is erased by the process $\rho_{S\!E}(t_{1})\rightarrow\rho_{S}(t_{1})\otimes\rho_{E}(t_{1})$ where $t_{1}=t_{I}$. Therefore, $\rho_{S}(t_{2})\neq\rho^{\prime}_{S}(t_{2})$ signifies that the environment (as well as its correlations with the system [39]) remembers the system’s history which in turn influences the system’s future.
The non-Markovnianity of is defined as [35]
@@
where $D(\rho_{1},\rho_{2})=\frac{1}{2}\|\rho_{1}-\rho_{2}\|$ is the trace distance with $\|A\|=\mathrm{Tr}(\sqrt{A^{\dagger}A})$ the trace norm of an operator $A$, and $\rho_{\Lambda}$ represents the Choi-Jamiółkowski matrices [42, 43] of a dynamical map $\Lambda$. It is given by $\rho_{\Lambda}=(\mathbb{I}\otimes\Lambda)(|\psi_{A\!S}\rangle\langle\psi_{A\!S}|)$ where $\mathbb{I}$ is the identity map and $|\psi_{A\!S}\rangle=\frac{1}{\sqrt{d}}\sum_{i=1}^{d}|i\rangle_{A}|i\rangle_{S}$ is a maximally entangled state between an ancillary system $A$ and the system $S$ (both $d$-dimensional). Using dynamical maps $T(t_{b},t_{a})$ in its definition, the measure does not depend on the system’s initial state. To investigate the influence of particular initial states on the memory effects, we recently use
@@
as the strength of memory effects conditioned on the initial state $\rho_{S}(t_{0})$ [36]. Note that $N_{M}[\rho_{S}(t_{0})]>0$ is a sufficient condition for the inequality Eq.(20) or $N_{M}>0$. In the following, the above two measures are both considered to deal with the control of both dynamical maps and quantum states by dynamical decoupling.